# AutoOPE: Automated Off-Policy Estimator Selection

### v1

@@ F.1 Scaling Experiment
We performed a scaling experiment of AutoOPE by varying the size of the meta-learning dataset $\mathcal{M}$. We take 20% of the OPE tasks in $\mathcal{M}$ as the test set. To investigate the scaling behavior, we train and optimize AutoOPE on different sub-samples of the remaining 80% of the OPE tasks in $\mathcal{M}$. The results are reported in Figure 5 and show that the loss (i.e., the Regret) of AutoOPE follows a scaling behavior, which is particularly popular in the recent Deep Learning literature [15, 11]. By increasing the training data, we observe a decrease in the Regret approximately following a power law in the form: $L=A+\frac{B}{D^{\alpha}}$, where $L$ is the loss measured with the Regret, $D$ is the meta-learning dataset size, and $A$, $B$, and $\alpha$ are scalar coefficients. While we selected the largest meta-learning dataset for the experimental analysis of this paper, the results in Figure 5 show that the dataset size can be adjusted based on the trade-off between computational resources and model performance for the scenario of interest.
Figure 5: Scaling experiment by varying the meta-learning dataset size.
F.2 Additional Ablation Studies
@@
We train 3 models (one for each group) in the same way outlined in Appendix D, and we test them on the same unseen test set used for the scaling experiment. We show the results in Table 5.
@@
F.3 Additional Experiments on CIFAR-10
The CIFAR-10 dataset [18] is a widely-used benchmark for image classification tasks, with mutually exclusive labeling.
CB Data
To adapt the CIFAR-10 dataset for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features (i.e., images pixels scaled in $(0,1)$) as context features and the class labels as actions. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using using parameters $\alpha_{b}$, $\alpha_{e}$:
[…]
Results
F.4 Additional Synthetic Experiments
In this section we describe all the experiments on synthetic data, following the experimental pipeline of [40].
CB Data
The synthetic logging data is generated using the following procedure and distributions (same procedure used in [40]):
  • A total of 10 possible actions is considered.
As already said, we divide the synthetic experiments into two main settings: one named Logging 1, where we set $\beta_{1}=2$ and $\beta_{2}=-2$, and the other called Logging 2, where $\beta_{1}=3$ and $\beta_{2}=7$, in line with the experimental analysis performed by Udagawa et al. [40].
To provide precise performance assessments, in all the 21 experiments of both Logging 1 and 2 we generate $n_{data}=100$ realizations of the logging dataset, as done in the procedure outlined in [40]. The performance metrics are subsequently averaged over all realizations to yield more robust performance estimates with their confidence intervals.
Results
The results obtained from Logging 1 are reported in Figure 8 while those of Logging 2 in Figure 8. Since these synthetic experiments are the same of [40], we use their optimal values for the other hyperparameters such as the network structure, regularization, learning rate. In both cases our proposed AutoOPE outperforms PAS-IF across almost all evaluation policies in terms of Spearman’s rank correlation coefficient. In terms of Relative Regret, AutoOPE performs similar to PAS-IF in the first half settings of the Logging 1 experiments. However, for high values of $\beta_{e}$, PAS-IF slightly outperform AutoOPE. In the Logging 2 set of experiments, we see that for all the $\beta_{e}\geq 0$ AutoOPE outperforms PAS-IF, but PAS-IF performs better in the settings with $\beta_{e}<0$. The fact that AutoOPE exhibits better Spearman’s Rank Correlation indicates that, while both methods predict similar values for the best estimator, AutoOPE is more consistent in ranking all of them.
Appendix G Infrastructure and Computational Time Comparison
@@
We used the c6i.32xlarge instance with 128 cores and 256 GB RAM for the synthetic dataset generation explained in Appendix C.2. This instance allows to terminate the whole generation process in about 35 hours.
For the AutoOPE model optimization, training and evaluation instead, additional RAM is needed, since some optimization iterations dive into the training of many and very deep decision trees. The instance used in this case is the m6i.32xlarge instance, with 128 cores and 512 GB RAM. The whole process last approximately 20 hours.
@@
Appendix H Ground-Truth Evaluation Policy Value Estimation
H.1 Open Bandit Dataset
@@
$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{gt}}\cdot\sum_{r_{t}\in\mathcal{D}_{gt}}r_{t}\right)^{2}$ (5)
H.2 UCI Datasets and CIFAR-10
In the experiment reported in Appendix F.3 and in the experiments in Section 4.2.2, we were able to compute the policy value of $\pi_{e}$ in the same way, since we use classification datasets. The policy value $\pi_{e}$ is computed as a true counterfactual expected reward on the logging dataset: in fact, the reward associated with each action and context in the logging data is deterministic (the reward is 1 for the correct class label/action relative to the context, and 0 for all other labels/actions). In this experiment the evaluation dataset can be defined as $\mathcal{D}_{gt}\coloneqq\left\{x_{t},\left\{r_{t,v}\right\}_{v}^{|\mathcal{A}|},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$. The estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with CIFAR-10 and UCI data (transformed in Contextual Bandit data as described in Appendix F.3 and in Section 4.2.2), is computed as described by the following expression:
$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{b}}\cdot\sum_{t}^{n_{b}}\sum_{v}^{|\mathcal{A}|}\pi_{e}(a_{v}|x_{t})\cdot r_{t,v}\right)^{2}$ (6)
H.3 Synthetic Data Experiments
@@
5:  $MSE(\hat{V}_{m}(\pi_{e}))\leftarrow\frac{1}{n_{e}}\cdot\sum_{s}^{n_{e}}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(s)})-\frac{1}{n_{gt}}\cdot\sum_{t}^{n_{gt}}\sum_{a_{j}\in\mathcal{A}}f_{\pi_{e}}(x_{t},a_{j};s)\cdot q(x_{t},a_{j};s))^{2}$
In Appendix F.4 are specified mostly of the input parameters for Logging 1 and 2 used in Algorithm 4, a part from $n_{e},\ n_{gt}$, set with values $n_{e}=100,\ n_{gt}=1,000,000$.
Appendix I Broader impact statement

### v2

@@ H.1 Scaling Experiment
We performed a scaling experiment of AutoOPE by varying the size of the dataset $\mathcal{M}$. We take 20% of the OPE tasks in $\mathcal{M}$ as the test set. To investigate the scaling behavior, we train and optimize AutoOPE on different sub-samples of the remaining 80% of the OPE tasks in $\mathcal{M}$. The results are reported in Figure 8 and show that the loss (i.e., the Regret) of AutoOPE follows a scaling behavior, which is particularly popular in the recent Deep Learning literature (Kaplan et al., 2020; Hoffmann et al., 2022). By increasing the training data, we observe a decrease in the Regret approximately following a power law in the form: $L=A+\frac{B}{D^{\alpha}}$, where $L$ is the loss measured with the Regret, $D$ is the meta-dataset size, and $A$, $B$, and $\alpha$ are scalar coefficients. While we selected the largest meta-dataset for the experimental analysis of this paper, the results in Figure 8 show that the dataset size can be adjusted based on the trade-off between computational resources and model performance for the scenario of interest.
Figure 8: Scaling experiment by varying the meta-dataset size.
H.2 Additional Ablation Studies
@@
We train 3 models (one for each group) in the same way outlined in Appendix E, and we test them on the same unseen test set used for the scaling experiment. We show the results in Table 5.
@@
Appendix I INFRASTRUCTURE AND COMPUTATIONAL TIME COMPARISON
@@
We used the c6i.32xlarge instance with 128 cores and 256 GB RAM for the synthetic dataset generation explained in Appendix D.2. This instance allows to terminate the whole generation process in about 35 hours.
For the AutoOPE model optimization, training and evaluation instead, additional RAM is needed, since some optimization iterations dive into the training of many and very deep decision trees. The instance used in this case is the m6i.32xlarge instance, with 128 cores and 512 GB RAM. The whole process last approximately 5 days.
@@
Appendix J GROUND-TRUTH EVALUATION AND POLICY VALUE ESTIMATION
J.1 Open Bandit Dataset
@@
$\text{MSE}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{gt}}\cdot\sum_{r_{t}\in\mathcal{D}_{gt}}r_{t}\right)^{2}$ (6)
J.2 UCI Datasets and CIFAR-10
In the experiment reported in Appendix G.1 and in the experiments in Section 4.2.2, we were able to compute the policy value of $\pi_{e}$ in the same way, since we use classification datasets. The policy value $\pi_{e}$ is computed as a true counterfactual expected reward on the logging dataset: in fact, the reward associated with each action and context in the logging data is deterministic (the reward is 1 for the correct class label/action relative to the context, and 0 for all other labels/actions). In this experiment the evaluation dataset can be defined as $\mathcal{D}_{gt}\coloneqq\left\{x_{t},\left\{r_{t,v}\right\}_{v}^{|\mathcal{A}|},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$. The estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with CIFAR-10 and UCI data (transformed in Contextual Bandit data as described in Appendix G.1 and in Section 4.2.2), is computed as described by the following expression:
$\text{MSE}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{b}}\cdot\sum_{t}^{n_{b}}\sum_{v}^{|\mathcal{A}|}\pi_{e}(a_{v}|x_{t})\cdot r_{t,v}\right)^{2}$ (7)
J.3 Synthetic Data Experiments
@@
5:  $\text{MSE}(\hat{V}_{m}(\pi_{e}))\leftarrow\frac{1}{n_{e}}\cdot\sum_{s}^{n_{e}}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(s)})-\frac{1}{n_{gt}}\cdot\sum_{t}^{n_{gt}}\sum_{a_{j}\in\mathcal{A}}f_{\pi_{e}}(x_{t},a_{j};s)\cdot q(x_{t},a_{j};s))^{2}$
In Appendix G.2 are specified mostly of the input parameters for Logging 1 and 2 used in Algorithm 4, a part from $n_{e},\ n_{gt}$, set with values $n_{e}=100,\ n_{gt}=1,000,000$.